# Reimagining Memory Access for LLM Inference: Compression-Aware Memory Controller Design

### v1

@@ Abstract
The efficiency of Large Language Model (LLM) inference is often constrained by substantial memory bandwidth and capacity demands. Existing techniques, such as pruning, quantization, and mixture of experts/depth, reduce memory capacity and/or bandwidth consumption at the cost of slight degradation in inference quality. This paper introduces a design solution that further alleviates memory bottlenecks by enhancing the on-chip memory controller in AI accelerators to achieve two main objectives: (1) significantly reducing memory capacity and bandwidth usage through lossless block compression (e.g., LZ4 and ZSTD) of model weights and key-value (KV) cache without compromising inference quality, and (2) enabling memory bandwidth and energy consumption to scale proportionally with context-dependent dynamic quantization. These goals are accomplished by equipping the on-chip memory controller with mechanisms to improve fine-grained bit-level accessibility and compressibility of weights and KV cache through LLM-aware configuration of in-memory placement and representation. Experimental results on publicly available LLMs demonstrate the effectiveness of this approach, showing memory footprint reductions of 25.2% for model weights and 46.9% for KV cache. In addition, our hardware prototype at 4 GHz and 32 lanes (7 nm) achieves 8 TB/s throughput with a modest area overhead (under 3.8 mm2), which underscores the viability of LLM-aware memory control as a key to efficient large-scale inference.
@@ I Introduction
To ensure practical feasibility, we propose integrating these techniques, along with hardware-accelerated lossless compression and decompression, into the on-chip memory controller of AI accelerators. This approach minimizes changes to the computational framework and software stack, as the memory controller only needs to be aware of the data semantics of weights and KV cache. All bit-level data manipulations and compression tasks occur within the memory controller, remaining transparent to the computing fabric and software stack. Through extensive experiments on publicly available LLMs, we demonstrate that our approach achieves significant reductions in memory footprint without compromising inference accuracy. Experimental results reveal memory footprint reductions of 25.2% for model weights and 46.9% for KV cache when using ZSTD lossless compression. Moreover, simulations conducted using DRAMSim3 [14] demonstrate that, compared to straightforward in-memory placement, our proposed approach can reduce data load latency by up to 32.0% and lower memory access energy consumption by 29.9%.
@@ II-C Dynamic Quantization
Dynamic quantization also extends to model weights, where precision adjustments based on contextual relevance allow finer granularity than dynamic pruning. Unlike traditional pruning methods, which only activate or deactivate model weights, dynamic quantization allows real-time precision adjustments (e.g., FP16, FP8, FP4), refining memory and energy efficiency by tuning precision to match computational needs. Within the Mixture of Depths and Experts (MoDE) framework [4], we implemented a context-dependent dynamic weight quantization, where routers control the precision level for each component of the model block, as illustrated in Fig. 2. A performance comparison on the LLaMA-MoE-3.5B model (Fig. 3) demonstrates the effectiveness of dynamic quantization in balancing memory efficiency and task accuracy across several configurations. For instance, a configuration quantizing more experts to lower precisions rather than simply skipping those experts achieved an improvement in zero-shot accuracy on the PIQA task, with a 1.9 percentage point increase over the baseline configuration. These results validate dynamic quantization as a robust method to enhance memory efficiency while sustaining performance across varied precision settings.
@@ IV-A1 KV Cache Compressibility
We evaluated the compressibility of the KV cache across 32 layers of the LLaMA 3.1 8B model on the WikiText dataset [26] and BookSum (LongBench) datasets [27]. In this work, we define compression ratio as ${S_{orig}}/{S_{comp}}\geq 1$, where $S_{orig}$ and $S_{comp}$ denote the size of the original and compressed data blocks. Compression ratios were measured for both LZ4 and ZSTD algorithms, as shown in Fig.7. Our data placement strategy on the KV cache achieved up to 46.9% overall footprint reduction. On WikiText, the highest compression ratios on a single layer reached 2.69 (ZSTD) and 2.31 (LZ4), while on BookSum, they peaked at 2.10 (ZSTD) and 1.93 (LZ4). Compared to the baseline (with overall ZSTD compression ratios of 1.21 on WikiText and 1.33 on BookSum), which does not apply cross-token KV cache clustering and de-correlation, our approach (with overall ratios of 1.81 on WikiText and 1.88 on BookSum) improves the overall KV cache lossless compression ratio by 50.3% on WikiText and 41.7% on BookSum using ZSTD.
Figure 7: Comparison of KV cache bit-plane 4KB compression ratios in the LLaMA 3.1 8B model (32 layers) using WikiText (test) datasets and BookSum task (LongBench). We evaluate cross-token KV cache clustering and de-correlation versus baseline approach with LZ4 and ZSTD compression.
@@ IV-A2 Model Weights Compressibility
Table III presents compression ratios and corresponding memory savings for various LLM configurations across different precision levels. BF16-based models achieve the highest lossless compression gains; for instance, LLaMA 3.1 8B in BF16 precision achieves a ZSTD lossless compression ratio of 1.34, leading to a 25.2% reduction. Moreover, since the proposed method is orthogonal to recent lossy compression techniques, it can also cooperate effectively with quantization approaches (e.g., GPTQ) to amplify total memory savings. For example, for FP8 and INT4 precision models, combining our method with AutoFP8 and GPTQ lossy compression—starting from BF16—results in substantial overall savings. For instance, quantizing LLaMA 3.1 8B to FP8 achieves a 54.1% total reduction, merging a 50% lossy saving with an additional 8.3% from our lossless compression approach.
TABLE III: Lossless compression ratios and total memory savings with lossy compression.
@@
Fig. 8 illustrates the compressibility of model weights across bit-planes for BF16, FP8, and INT4-based LLMs using 4KB ZSTD lossless compression, along with KV cache compressibility in the LLaMA 3.1 8B model on WikiText and BookSum datasets. For BF16 model weights, the top four exponent bit-planes contribute the most to overall compressibility, achieving an overall compression ratio of 1.34. This is because exponents, especially in high-precision formats like BF16, often contain more redundancy and fewer unique values, allowing for effective compression. In contrast, FP8 and INT4 models, already subjected to lossy quantization, show limited compressibility as the reduced bit precision minimizes representational redundancy, particularly in the exponent bits.
@@
For the KV cache in BF16 format, shown in the lower two subfigures of Fig. 8, the exponent bit-planes again demonstrate significantly higher compressibility. This is attributed to the relatively narrow range of data stored along the channel in KV cache, where exponents frequently exhibit low variability across tokens. These properties lead to substantial memory savings of 44.8% on the WikiText and 46.9% on BookSum.
Figure 9: Precision distribution for models LLaMA 3.1 8B, LLaMA 3.1 70B, Mixtral 8×7B, and LLaMA-MoE-3.5B in BF16, quantized FP8, and INT4 when conducting inference on WikiText-2.
IV-B DRAM Access Efficiency
@@
Fig. 9 shows 12 LLMs in aforementioned configuration and their precision distribution when conducting inference on WikiText-2 [26]. The inference were configured with various quantization schemes: BF16, FP12/8/6/4 applied to BF16-based models; FP8/6/4 applied to FP8-based models; and INT4/2 applied to INT4-based models. All router layers are using BF16 precision for accuracy.
Fig. 10 compares the DRAM access energy, detailing activation and read components, between the proposed bit-plane method (P) and a traditional straightforward byte-level approach (T). Implementing the proposed method results in a DRAM access energy reduction of up to 29.9%. Specifically, for BF16-based models, the proposed method achieves energy reductions of 27.8% for LLaMA 3.1 8B, 25.9% for LLaMA 3.1 70B, 29.9% for Mixtral 8×7B, and 27.2% for LLaMA MoE 3.5B. In the case of the Mixtral 8×7B model, the proposed method reduces DRAM access energy by 19.6% for the FP8-based model and 17.9% for the INT4-based model compared to the traditional approach, which shows a decreasing trend in energy savings as the quantization precision of the model decreases.
@@
Fig. 11 shows the average model load latency for aforementioned 12 LLMs during inference on the WikiText-2 dataset. The proposed method achieves latency reductions of up to 32.0%. Specifically, for BF16-based models, the proposed method reduces load latency by 29.8% for LLaMA 3.1 8B, 26.5% for LLaMA 3.1 70B, 32.0% for Mixtral 8×7B, and 26.5% for LLaMA-MoE-3.5B.
@@ IV-C Hardware Implementation and Resource Evaluation
Table IV summarizes the hardware resource usage at 4 GHz with 32 parallel lanes for both LZ4 and ZSTD compression engines (7nm PDKs), evaluated over three block sizes (16 K, 32 K, and 64 K bits). As the block size increases, we observe that Single-Lane Area (SL Area) for LZ4 rises from 0.038 mm2 at 16 K bits to 0.101 mm2 at 64 K bits, roughly a 2.7$\times$ increase. A similar trend occurs for ZSTD, ranging from 0.056 mm2 up to 0.119 mm2. Single-Lane Power (SL Power) shows a proportional escalation, e.g., LZ4’s single-lane power goes from 929 mW to 2187 mW. ZSTD requires more complex entropy coding, leading to a consistently higher single-lane power (up to 3077 mW). Lane-Total Area (LaneTotArea) aggregates the single-lane area for 32 lanes. At 64 K bits, LZ4 exhibits 3.223 mm2 while ZSTD reaches 3.796 mm2. Each lane delivers 2048 Gbps (256 GB/s). With 32 lanes, the system sustains 65,536 Gbps (8 TB/s), comfortably surpassing a nominal 2 TB/s memory interface target.
TABLE IV: Parameter sweep results at 4 GHz with 32 lanes for LZ4 and ZSTD compression. (“SL” denotes single-lane.)
Engine BlockSize SL Area SL Power LaneTotArea LaneTotPower SL Thpt
(bits) (mm2) (mW) (mm2) (mW) (Gbps)
LZ4 16384 0.03779 928.686 1.20942 29717.955 2048
LZ4 32768 0.05038 1180.344 1.61207 37771.018 2048
LZ4 65536 0.10071 2186.977 3.22268 69983.273 2048
ZSTD 16384 0.05571 1818.286 1.78286 58185.155 2048
ZSTD 32768 0.06830 2069.944 2.18551 66238.218 2048
ZSTD 65536 0.11863 3076.577 3.79612 98450.473 2048
@@ VI Conclusion
This paper has presented a memory controller–centric solution for mitigating the bandwidth and capacity bottlenecks that commonly arise in LLM inference. By judiciously reorganizing floating-point data at the bit level (bit-plane disaggregation) and exploiting cross-token correlation in the KV cache, our design enables efficient application of standard lossless compressors (LZ4, ZSTD). Experimental results on a range of publicly available LLMs demonstrate that this approach reduces model weights by up to 25.2% and KV cache by as much as 46.9%—all without degrading model accuracy. Moreover, the proposed layout allows dynamic quantization to scale memory bandwidth and energy consumption proportionally, achieving up to 32.0% faster model load times and 29.9% lower DRAM access energy in DRAMSim3 simulations. Our hardware evaluation further reveals that even when instantiated at 4 GHz across 32 parallel lanes (7 nm PDK), the total area overhead remains modest (e.g., 3.22 mm2 for an LZ4-based design and 3.80 mm2 for ZSTD), while delivering 8 TB/s of effective throughput. These findings underscore the pivotal role that an LLM-aware memory controller can play in meeting the growing performance and efficiency demands of modern AI inference workloads.

### v3

@@ Abstract
The efficiency of Large Language Model (LLM) inference is often constrained by substantial memory bandwidth and capacity demands. Existing techniques, such as pruning, quantization, and mixture of experts/depth, reduce memory capacity and/or bandwidth consumption at the cost of slight degradation in inference quality. This paper introduces a design solution that further alleviates memory bottlenecks by enhancing the on-chip memory controller in AI accelerators to achieve two main objectives: (1) significantly reducing memory capacity and bandwidth usage through lossless block compression (e.g., LZ4 and ZSTD) of model weights and key-value (KV) cache without compromising inference quality, and (2) enabling memory bandwidth and energy consumption to scale proportionally with context-dependent dynamic quantization. These goals are accomplished by equipping the on-chip memory controller with mechanisms to improve fine-grained bit-level accessibility and compressibility of weights and KV cache through LLM-aware configuration of in-memory placement and representation. Experimental results on publicly available LLMs demonstrate the effectiveness of this approach, showing memory footprint reductions of 25.2% for model weights and 46.9% for KV cache. In addition, our hardware simulation results at 2 GHz and 32 lanes achieve up to 2 TB/s throughput with a modest area overhead (under 5.69 mm2), which underscores the viability of LLM-aware memory control as a key to efficient large-scale inference.
@@ I Introduction
To ensure practical feasibility, we propose integrating these techniques, along with hardware-accelerated lossless compression and decompression, into the on-chip memory controller of AI accelerators. This approach minimizes changes to the computational framework and software stack, as the memory controller only needs to be aware of the data semantics of weights and KV cache. All bit-level data manipulations and compression tasks occur within the memory controller, remaining transparent to the computing fabric and software stack. Through extensive experiments on publicly available LLMs, we demonstrate that our approach achieves significant reductions in memory footprint without compromising inference accuracy. Experimental results reveal memory footprint reductions of 25.2% for model weights and 46.9% for KV cache when using ZSTD lossless compression. Moreover, simulations conducted using DRAMSim3 [14] demonstrate that, compared to straightforward in-memory placement, our proposed approach can reduce data load latency by up to 30.0% and lower memory access energy consumption by 29.9%.
@@ II-C Dynamic Quantization
Dynamic quantization also extends to model weights, where precision adjustments based on contextual relevance allow finer granularity than dynamic pruning. Unlike traditional pruning methods, which only activate or deactivate model weights, dynamic quantization allows real-time precision adjustments (e.g., FP16, FP8, FP4), refining memory and energy efficiency by tuning precision to match computational needs. Within the Mixture of Depths and Experts (MoDE) framework [4], we implemented a context-dependent dynamic weight quantization, where routers control the precision level for each component of the model block, as illustrated in Fig. 2. A performance comparison on the LLaMA-MoE-3.5B model (Fig. 3) demonstrates the effectiveness of dynamic quantization in balancing memory efficiency and task accuracy across several configurations. For instance, a configuration quantizing more experts to lower precisions rather than simply skipping those experts achieved an improvement in zero-shot accuracy on the PIQA task, with a 1.9 percentage points increase over the baseline configuration. These results validate dynamic quantization as a robust method to enhance memory efficiency while sustaining performance across varied precision settings.
@@ IV-A1 KV Cache Compressibility
We evaluated the compressibility of the KV cache across 32 layers of the LLaMA 3.1 8B model on the WikiText dataset [26] and BookSum (LongBench) datasets [27]. In this work, we define compression ratio as ${S_{orig}}/{S_{comp}}\geq 1$, where $S_{orig}$ and $S_{comp}$ denote the size of the original and compressed data blocks. Compression ratios were measured for both LZ4 and ZSTD algorithms with a 4KB compression block size, as shown in Fig.7. Our data placement strategy on the KV cache achieved 44.8% and 46.9% overall footprint reduction on WikiText and Booksum task. On WikiText, the highest compression ratios on a single layer reached 2.69 (ZSTD) and 2.31 (LZ4), while on BookSum, they peaked at 2.10 (ZSTD) and 1.93 (LZ4). Compared to the baseline (with overall ZSTD compression ratios of 1.21 on WikiText and 1.33 on BookSum), which does not apply cross-token KV cache clustering and de-correlation, our approach (with overall ratios of 1.81 on WikiText and 1.88 on BookSum) improves the overall KV cache lossless compression ratio by 50.3% on WikiText and 41.7% on BookSum using ZSTD.
Figure 7: Comparison of KV cache bit-plane 4KB compression ratios in the LLaMA 3.1 8B model (32 layers) using WikiText (test) datasets and BookSum task (LongBench). We evaluate cross-token KV cache clustering and de-correlation versus baseline approach with LZ4 and ZSTD compression with 4KB compression block.
@@ IV-A2 Model Weights Compressibility
Table III presents compression ratios and corresponding memory savings for various LLM configurations across different precision levels on model weights. BF16-based models achieve the highest lossless compression gains; for instance, LLaMA 3.1 8B in BF16 precision achieves a ZSTD lossless compression ratio of 1.34, leading to a 25.2% reduction. Moreover, since the proposed method is orthogonal to recent lossy compression techniques, it can also cooperate effectively with quantization approaches (e.g., GPTQ [3]) to amplify total memory savings. For example, for FP8 and INT4 precision models, combining our method with AutoFP8 and GPTQ lossy compression—starting from BF16—results in substantial overall savings. For instance, quantizing LLaMA 3.1 8B to FP8 achieves a 54.1% total reduction, merging a 50% lossy saving with an additional 8.3% from our lossless compression approach.
TABLE III: Lossless compression ratios and total memory savings with lossy compression on model weights
@@
Fig. 8 illustrates the compressibility of model weights across bit-planes for BF16, FP8, and INT4-based LLMs using 4KB ZSTD lossless compression, along with KV cache compressibility in the LLaMA 3.1 8B model on WikiText and BookSum datasets. For BF16 model weights, the top four exponent bit-planes contribute the most to overall compressibility, and together with other bit-planes it achieves an overall compression ratio of 1.34. This is because exponents, especially in high-precision formats like BF16, often contain more redundancy and fewer unique values, allowing for effective compression. In contrast, FP8 and INT4 models, already subjected to lossy quantization, show limited compressibility as the reduced bit precision minimizes representational redundancy, particularly in the exponent bits.
@@
For the KV cache in BF16 format, shown in the lower two subfigures of Fig. 8, the exponent bit-planes again demonstrate significantly higher compressibility. This is attributed to the relatively narrow range of data stored along the channel in KV cache, where exponents frequently exhibit low variability across tokens. These properties lead to substantial memory savings of 44.8% on the WikiText and 46.9% on BookSum (as shown in Fig. 7).
Figure 9: Precision distribution for model weights of LLaMA 3.1 8B, LLaMA 3.1 70B, Mixtral 8×7B, and LLaMA-MoE-3.5B in BF16, quantized FP8, and INT4 when conducting inference on WikiText-2.
IV-B DRAM Access Efficiency with Dynamic Quantization
@@
Fig. 9 shows 12 LLMs in aforementioned configuration and their average model weights precision distribution when conducting inference on WikiText-2 [26]. The inference were configured with various dynamic quantization schemes: BF16, FP12/8/6/4 applied to BF16-based models; FP8/6/4 applied to FP8-based models; and INT4/2 applied to INT4-based models. All router layers are using BF16 precision for accuracy.
Fig. 10 compares the DRAM access energy with dynamic quantization, and shows read and activation energy between the proposed bit-plane method (P) and a traditional straightforward byte-level approach (T). Implementing the proposed method results in a DRAM access energy reduction of up to 29.9%. Specifically, for BF16-based models, the proposed method achieves energy reductions of 27.8% for LLaMA 3.1 8B, 25.9% for LLaMA 3.1 70B, 29.9% for Mixtral 8×7B, and 27.2% for LLaMA MoE 3.5B. In the case of the Mixtral 8×7B model, the proposed method reduces DRAM access energy by 19.6% for the FP8-based model and 17.9% for the INT4-based model compared to the traditional approach, which shows a decreasing trend in energy savings as the quantization precision of the model decreases.
@@
Fig. 11 presents the average model load latency on the WikiText-2 dataset for aforementioned 12 LLMs when applying same dynamic quantization. The proposed method achieves latency reductions of up to 30.0%. In Mixtral 8×7B, the BF16 configuration shows a decrease from 705.90 to 495.06 ms (a 30.0% reduction), while the FP8 and INT4 formats achieve respective reductions of approximately 17.1% and 14.5%. For the larger LLaMA 3.1 70B model, BF16 latency is cut from 910.58 to 674.73 ms (25.9% reduction), with the FP8 and INT4 versions also exhibiting notable improvements (from 348.65 to 293.27 ms and from 251.03 to 214.11 ms, respectively). These results demonstrate that the proposed design not only minimizes DRAM traffic but also significantly reduces load latency across various architectures and precision levels, thereby boosting overall inference efficiency.
@@
TABLE IV: Silicon cost at 2 GHz with 32 lanes for LZ4 and ZSTD lossless compression. (“SL” denotes single-lane.)
Engine BlockSize SL Area SL Power LaneTotArea LaneTotPower SL Thpt
(bits) (mm2) (mW) (mm2) (mW) (Gbps)
LZ4 16384 0.05669 696.515 1.81413 2228.846 512
LZ4 32768 0.07557 885.258 2.41811 2832.826 512
LZ4 65536 0.15106 1640.233 4.83403 5248.745 512
ZSTD 16384 0.08357 1363.715 2.67429 4363.886 512
ZSTD 32768 0.10245 1552.458 3.27827 4967.866 512
ZSTD 65536 0.17794 2307.433 5.69419 7384.785 512
@@ IV-C Hardware Implementation and Resource Evaluation
Table IV summarizes the hardware resource usage at 2 GHz with 32 parallel lanes for both LZ4 and ZSTD compression engines (synthesized with 7 nm PDKs [33]), evaluated over three block sizes (16384, 32768, and 65536 bits). For LZ4, the single lane area increases from 0.057 mm2 at 16384 bits to 0.151 mm2 at 65536 bits, while the Single-Lane Power increases from 696.515 mW to 1640.233 mW. For ZSTD, the Single-Lane Area increases from 0.084 mm2 to 0.178 mm2, and the Single-Lane Power rises from 1363.715 mW to 2307.433 mW. Aggregating across 32 lanes, LZ4 achieves total areas of 1.814 mm2 (16384 bits), 2.418 mm2 (32768 bits), and 4.834 mm2 (65536 bits), with corresponding total powers of 2228.846 mW, 2832.826 mW, and 5248.745 mW. Similarly, ZSTD achieves total areas of 2.674 mm2, 3.278 mm2, and 5.694 mm2 with total powers of 4363.886 mW, 4967.866 mW, and 7384.785 mW, respectively. Each lane delivers a throughput of 512 Gbps, resulting in an aggregate throughput of 16384 Gbps (2 TB/s) for 32 lanes.
@@ VI Conclusion
This paper has presented a memory controller–centric solution for mitigating the bandwidth and capacity bottlenecks that commonly arise in LLM inference. By judiciously reorganizing floating-point data at the bit level and exploiting cross-token correlation in the KV cache, our design enables efficient application of standard lossless compressors (LZ4, ZSTD). Experimental results on a range of publicly available LLMs demonstrate that this approach reduces model weights by up to 25.2% and KV cache by as much as 46.9%—all without degrading model accuracy. Moreover, the proposed layout allows dynamic quantization to scale memory bandwidth and energy consumption proportionally, achieving up to 30.0% faster model load times and 29.9% lower DRAM access energy in DRAMSim3 simulations. Our hardware evaluation further reveals that even when instantiated at 2 GHz across 32 parallel lanes, the total area overhead remains modest (e.g., 4.83 mm2 for an LZ4-based design and 5.69 mm2 for ZSTD), while delivering up to 2 TB/s of effective throughput. These findings underscore the pivotal role that an LLM-aware memory controller can play in meeting the growing performance and efficiency demands of modern AI inference workloads.